OpenStudy (anonymous):

5,8,10,20,25,30,35,40,45 what is the lower quartile of data set above. A.5 B.5 C.9 D.10

OpenStudy (texaschic101):

5,8,10,20,25,30,35,40,45 Quartile 1 is your lower quartile Q2 is your median...middle number Q3 is your upper quartile Q2 would be 25 Q1 (upper quartile) is the median of the numbers before Q2 so Q1 is : (8 + 10) / 2 = 18/2 = 9 (your lower quartile).